Underarmour forecasts sales of $5 Billion in 2017. They are trying to decide how much to allocate to advertising and their agency has told them that their competitors in the athletic apparel/accessories segment spend on average 3.2% of sales on advertising. They spent $140 million on advertising in 2016. Using the A/S ratio method, calculate a budget for Underarmour for 2017:

a.

$16,000,000

b.

$140,000,000

c.

$150,000,000

d.

$160,000,000

e.

None of these
Answer #1

Sales (S) = 5 Billion in 2017

A/S = 3.2% (the competition is th benchmark)

So, Advertisement budget in 2017 (A) = S * 3.2% = 5 Billion * 3.2% = 0.16 billion i.e. $160,000,000